# A Continuous Energy Ising Machine Leveraging Difference-of-Convex Programming

### v1

@@ Abstract
Many combinatorial optimization problems can be reformulated as the task of finding the ground state of a physical system, such as the Ising model. Most existing Ising solvers are inspired by simulated annealing. Although annealing techniques offer scalability, they lack convergence guarantees and are sensitive to the cooling schedule. We propose to solve the Ising problem by relaxing the binary spins to continuous variables and introducing a potential function (attractor) that steers the solution toward binary spin configurations. The resulting Hamiltonian can be expressed as a difference of convex functions, enabling the design of efficient iterative algorithms that require a single matrix-vector multiplication per iteration and are backed by convergence guarantees. We implement our Ising solver across a range of GPU platforms—from edge devices to high-performance computing clusters—and demonstrate that it consistently outperforms existing solvers across problem sizes ranging from small ($10^{3}$ spins) to ultra-large ($10^{8}$ spins).
@@ 1 Introduction
Combinatorial optimization problems appears in numerous fields such as social networks [1], finance [2], cryptography [3], scheduling [4], electronic circuit design [5], and biosciences [6, 7]. The challenge with such problems is that the number of possible solutions grows exponentially with the number of variables, rendering classical algorithms slow or impractical for large-scale problems. To address this, researchers are increasingly exploring unconventional computing paradigms. A particularly promising direction is to reformulate the problem as finding the ground state of a physical system, such as the Ising model. Indeed, many Nondeterministic Polynomial-time (NP)-hard problems can be naturally expressed in the Ising form, while many others can be efficiently reduced to this framework [8]. Significant effort has been directed toward developing specialized hardware and algorithms, collectively known as Ising machines, that can efficiently compute the ground states of Ising models [9]. As industrial optimization problems grow in complexity, building large-scale Ising solvers has become more critical than ever. Beyond its role in combinatorial optimization, finding the ground state of the Ising model is a fundamental problem in physics, offering insights into phase transitions, magnetism, and critical behaviour in condensed matter systems.
The Ising model was originally introduced by Lenz and Ising as a mathematical model for ferromagnetism [10]. For an Ising model, the dimensionless energy function is given by
@@
where the spins $s_{1},\ldots,s_{n}$ take $\{-1,+1\}$ values (up/down spins), $J_{ij}$ is the coupling between spins $s_{i}$ and $s_{j}$ and $J_{ii}=0$ (no self-coupling), and $h_{i}$ is the external field acting on spin $s_{i}$. Solving an Ising model refers to finding a spin assignment $s^{*}_{1},\dots,s^{*}_{n}\in\{-1,1\}$ that minimizes the energy over all possible spin configurations, that is,
@@
We will refer to $(s^{*}_{1},\dots,s^{*}_{n})$ as the ground state spin vector (or simply the ground state) of the Ising model.
@@
Solving large Ising models by brute force is computationally challenging, with no known polynomial-time algorithm for the general case. In fact, the problem is known to be NP-hard [11]. Consequently, rather than attempting to compute the exact ground state, practical approaches typically focus on finding good approximations. Over the years, many algorithmic heuristics and specialized hardware have been developed to address this challenge. We refer the reader to [12, 13, 14, 15] for a comprehensive survey of past and recent advances.
@@
The most widely used technique for solving Ising models is simulated annealing (SA) [16, 17, 18, 19], a method inspired by the principles of quantum annealing [20, 21, 22]. SA uses Monte Carlo sampling guided by the Boltzmann distribution to search for the ground state. However, it relies on heuristic cooling schedules to explore the energy landscape and does not come with any mathematical guarantees. Moreover, the need for gradual temperature reduction can result in slow convergence [23], making SA less suited for solving large-scale problems. Variants such as noisy mean field annealing (NMFA) [24, 25] and mean field annealing from a random state (MARS) [23] aim to speed up convergence. However, they are typically prone to approximation errors and sensitive to hyperparameter settings. State-of-the-art techniques based on Hamiltonian dynamics, such as Simulated Bifurcation Machine (SBM) [26, 27, 28, 29, 30] and its ballistic variant (bSB) [27], offer parallelism but remain sensitive to parameter tuning and cooling schedules. On the other hand, Coherent Ising Machines (CIMs) [31, 32, 33], which leverage optical hardware, suffer from issues like noise, decoherence, and the need for complex, precisely controlled hardware [34, 35]. More recently, neural network-based approaches have shown promise [36, 37, 38]; however, they face inherent challenges, including the need for careful hyperparameter tuning and the difficulty in scaling to the vast configuration of large problems [38]. Thus, there is a need to develop Ising solvers that can deliver both high-speed performance and reliable solution quality, particularly for ultra-large-scale problems.
@@
Figure 1: Overview of our Ising solver for a $2$-spin system. For a $2$-spin Ising model, we begin by relaxing the $2^{2}=4$ discrete energy values $(\mathcal{E}_{0},\mathcal{E}_{1},\mathcal{E}_{2},\mathcal{E}_{3})$ to obtain a smooth energy landscape $\mathcal{E}(x_{1},x_{2})$ over the continuous space. We augment this energy with an attractor $\mathcal{A}(x_{1},x_{2})$––with tunable parameters $\alpha$ and $\beta$––to form the Hamiltonian $\mathcal{H}=\mathcal{E}+\mathcal{A}$. The local minima of this Hamiltonian, denoted by green dots $(\mathcal{H}_{0},\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3})$, correspond to the candidate ground states. By appropriately tuning $\alpha$ and $\beta$, we can ensure that the minimizer $(x^{*}_{1},x^{*}_{2})$ of the tuned Hamiltonian recovers the ground state of the Ising model i.e. $\mathrm{sign}(x^{*}_{1},x^{*}_{2})$ closely matches the true ground state $\boldsymbol{s}^{*}=(s^{*}_{1},s^{*}_{2})$ of the original Ising energy $\mathcal{E}$
@@
In this work, we propose a continuous optimization-based Ising solver. Unlike the classical Goemans-Williamson Semidefinite Program [39], our method avoids the computational bottleneck associated with large semidefinite programs (Supplementary Note 5). To improve the scalability, we drop the binary spin constraints entirely and instead couple a potential function (attractor) with the Ising energy, resulting in a Hamiltonian with soft spin constraints. A key insight is that this Hamiltonian can be expressed as a difference of convex functions, enabling the use of the powerful framework of difference-of-convex programming (DCP) [40, 41], which, to our knowledge, has not been applied to the Ising model. We present two DCP-based iterative solvers that require a single matrix-vector multiplication per iteration and have just two tunable parameters. By exploiting the mathematical properties of our Hamiltonian, we prove that one of our solvers is guaranteed to converge to a critical point, and under mild assumptions, to a local minimum, a property that is challenging to establish in nonconvex optimization. Our solvers are well-suited for parallel execution on advanced graphical processing units (GPU) and are versatile enough to run efficiently across a wide range of platforms, from low-power edge devices to high-performance computing clusters. We demonstrate that they consistently outperform existing Ising solvers across a broad range of problem sizes, including ultra-large-scale instances with up to $10^{8}$ spins. Notably, we can solve a fully connected $10^{7}$-spin Ising model with nearly $50$ trillion coupling terms in just $14$ hours using four NVIDIA H100 GPUs. A quick comparison with existing Ising solvers is provided in Table 1.
Annealing Markov Chain Monte Carlo (MCMC) SA [16, 17], MARS [23], NFMA [24] high yes poor no no hard
Simulated Hamiltonian Dynamics bSB [26, 27] high yes good no no hard
CIM [32, 33] high yes good no no hard
SIA [38] high yes good no no hard
Continuous Optimization CP [42] low no poor yes yes
SDP [43] low no poor yes yes
DOCH (present work) high yes good yes yes easy
Table 1: Comparison of key aspects of our Ising solver with existing solvers. SA: Simulated Annealing, MARS: Mean field Annealing from a Random State, NFMA: Noisy Mean Field Annealing, bSB: ballistic Bifurcation Machine, CIM: Coherent Ising Machine, SIA: Spring Ising Algorithm, CP: Convex Programming, SDP: Semidefinite Programming, DOCH: Difference of Convex Hamiltonian. (A convergence guarantee means that the iterative process will asymptotically reach a fixed point or a local minimum. In contrast, the approximation guarantee refers to how close the obtained solution is to the true global minimum (ground state) of the system).
@@ 2 Results
We develop a continuous energy model by embedding the spins in $\mathbb{R}^{n}$ (Figure 1). Consider the homogeneous Ising problem:
$\underset{\boldsymbol{s}\in\{-1,1\}^{n}}{\min}\ \mathcal{E}(\boldsymbol{s})=-\frac{1}{2}\boldsymbol{s}^{\top}\mathbf{J}\boldsymbol{s}$ (4)
where $\mathbf{J}=(J_{ij})$ is the symmetric coupling matrix ($J_{ii}=0$). We can identify the optimization space in (4) with the vertices of the unit hypercube in $\mathbb{R}^{n}$, and leverage the quadratic structure of $\mathcal{E}$ to expand the domain to $\{-\lambda,\lambda\}^{n}$ for any $\lambda>0$. More precisely, we consider the problem
@@
where $\mathbb{D}=\big{\{}\{-\lambda,\lambda\}^{n}:\ \lambda>0\big{\}}$ and $\boldsymbol{x}=(x_{1},\ldots,x_{n})$ is the optimization variable. The domain $\mathbb{D}$ consists of radial lines emanating from the origin along the vertices of $\{-1,1\}^{n}$. We can show that the optimization problems (4) and (5) are equivalent: if $\boldsymbol{x}^{*}$ is a global minimizer of (5), then $\boldsymbol{s}^{*}=\mathrm{sign}(\boldsymbol{x}^{*})$ is the ground state of (4).
Although formulation (5) brings us closer to the continuum, the problem is that $\mathbb{D}$ is a non-convex set that is not even connected. A natural idea is to consider its convex hull. However, this leads to the entire space $\mathbb{R}^{n}$, making the relaxation too loose and, therefore, ineffective for optimization. To address this, we introduce a potential function (attractor) that biases the minimizers towards $\mathbb{D}$. In particular, for fixed $\alpha,\beta>0$, we consider the attractor
@@
The parameters $\alpha$ and $\beta$ are used to control the shape of the attractor (Figure 2).
Figure 2: Shape of the attractor. (a) Plot of the attractor in one variable for different $\alpha,\beta$ values. The global minimizers at $\pm(\alpha/\beta)^{1/2}$ are shown. (b) Plot of the attractor function in two variables for $\alpha=\beta=1$, exhibiting four global minimizers at (1,1),(1,1),(1,1)(1,1),(1,-1),(-1,1) and (1,1)(-1,-1).
This specific form was motivated by the Hamiltonian of nonlinear parametric oscillators [44] (Supplementary Note 4). Although this provides a strong physical motivation, our primary interest in (6) lies in its mathematical properties. In particular, the global minimizers of $\mathcal{A}$ are exactly the vertices of the hypercube $\{-\lambda,\lambda\}^{n},\lambda=(\alpha/\beta)^{1/2}$ (Supplementary Note 2). Thus, by coupling $\mathcal{A}$ with the Ising energy, we bias the solutions towards $\mathbb{D}$. More specifically, we define the Hamiltonian to be $\mathcal{H}=\mathcal{A}+\mathcal{E}$, and consider the optimization problem
@@
Unlike the original Ising problem (4) that has a finite search space, a difficulty with continuous optimization problems is that they might not have a minimizer, i.e., the problem may not be well-posed. Nevertheless, we can prove that $\mathcal{H}$ is bounded below and always has a global minimizer $\boldsymbol{x}^{*}$ (Theorem S1 in Supplementary Note 2). Moreover, if it so happens that $\boldsymbol{x}^{*}\in\mathbb{D}$, then $\mathrm{sign}(\boldsymbol{x}^{*})$ is guaranteed to be the ground state of the original Ising problem (Supplementary Note 2).
@@
A direct approach would be to locate a critical point of $\mathcal{H}$ by setting its gradient to zero, which results in a system of cubic equations. However, solving nonlinear equations is a challenging task. While iterative methods such as gradient descent offer a more practical alternative, they require careful tuning of the step size and the parameters $\alpha$ and $\beta$. We propose an iterative algorithm that not only has a significantly lower per-iteration cost than gradient descent but also comes with a formal convergence guarantee. This is based on the observation that we can write the Hamiltonian $\mathcal{H}$ as the difference of convex functions for proper settings of $\alpha$ and $\beta$. More precisely, combining the quadratic components of $\mathcal{A}$ and $\mathcal{E}$, we have
@@
The function $f$ is convex for any $\beta>0$. On the other hand, $g$ is convex if and only if $\lambda_{\min}(\mathbf{J}+\alpha\mathbf{I})\geqslant 0$, where $\lambda_{\min}$ is the smallest eigenvalue. This can easily be guaranteed by making $\alpha$ sufficiently large (Supplementary Note 2). Under these conditions, the optimization problem (7) becomes
@@
where $f$ and $g$ are convex functions. This falls under the purview of difference-of-convex programming (DCP) [40, 41]. We consider a simple DCP algorithm called DCA [45], along with its accelerated variant [41].
DCA builds upon the classical principle of bound optimization [46]. Starting with an initialization $\boldsymbol{x}^{(0)}$, DCA generates a sequence of estimates $\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)},\cdots$ that is expected to converge to a minimizer of $\mathcal{H}$. The core idea is to exploit the convexity of $g$ to construct a global convex upper bound on fgf-g around the current estimate $\boldsymbol{x}^{(k)}$. This surrogate is then minimized to obtain the next iterate $\boldsymbol{x}^{(k+1)}$. Applied to problem (9), we obtain a simple update rule:
@@
where $\mathcal{T}$ is a linear transform followed by a pointwise nonlinearity (see Methods). In other words, the algorithm reduces to the repeated application of the operator $\mathcal{T}$, referred to as a fixed-point algorithm. The cost per iteration is just a matrix-vector multiplication. A simple example illustrating the behavior of this iterative algorithm is shown in Figure 3, using a toy Ising model to visualize the update steps and convergence.
@@
Figure 3: Hamiltonian and optimization trajectory for a $2$-spin system. We consider the anti-ferromagnetic model with $J_{12}=J_{21}=-1$, whose ground states are (1,1)(1,-1) and (1,1)(-1,1). Top row: Surface plots of the attractor $\mathcal{A}$, Ising energy $\mathcal{E}$, and the Hamiltonian $\mathcal{H}=\mathcal{A}+\mathcal{E}$. Bottom row: Corresponding contour plots for $\alpha=1$ and $\beta=2$. The white curves trace the optimization trajectories of our Ising solver, starting from various initial points (green dots). Depending on the initialization, the iterates converge to one of the two degenerate ground states (red dots) within five iterations.
We also consider a variant of DCA that achieves improved convergence rates using a technique called Nesterov acceleration [47]. We refer to this as Accelerated DCA (ADCA), which is described in detail in Methods. The solution quality achieved by ADCA consistently outperforms DCA as the problem size increases. However, while the ADCA iterates are found to converge in practice, it is difficult to provide a convergence guarantee.
@@
where $\eta$ is a tuning parameter. It can be shown that the smallest eigenvalue of $\mathbf{J}$ is negative, so that $\lambda_{\max}(-\mathbf{J})>0$ in (10). The above choice of $\alpha$ ensures that $\mathbf{J}+\alpha\mathbf{I}$ is positive semidefinite, guaranteeing the convexity of $g$ in (9).On the other hand, the choice of $\beta$ ensures boundedness of the DCA iterates for any value of $\alpha$ (Supplementary Note 3). We found that this particular setting consistently yields good results. For large models ($n\geqslant 10^{4}$), computing $\lambda_{\max}(\mathbf{J})$ can be computationally demanding. To address this, we approximate $\lambda_{\max}(\mathbf{J})$ using Wigner’s semicircle law [48], following the approach in [26, 27] for estimating eigenvalues of large matrices. Specifically, we estimate $\lambda_{\max}(\mathbf{J})$ with $2\langle\mathbf{J}\rangle\sqrt{n}$, where $\langle\mathbf{J}\rangle$ denotes the sample variance of the entries of $\mathbf{J}$:
@@
We tune the parameter $\eta$ in (10) by running a small number of iterations of our Ising solver, observing its behaviour, and adjusting $\alpha$ accordingly (Supplementary Note 6). This straightforward tuning procedure and the low per-iteration cost make our DCA-based solvers particularly well suited for large and ultra-large-scale Ising problems.
Our overall approach is summarized in Figure 1, starting with the relaxation of the binary spin constraints, incorporation of the attractor function with the Ising energy, DCP formulation of the Hamiltonian, tuning the parameters $\alpha,\beta$ and finally the use of DCA for optimizing the tuned Hamiltonian. We refer to this approach as DOCH (Difference-Of-Convex-Hamiltonian) and the accelerated variant as ADOCH.

### v2

@@ Abstract
Many combinatorial optimization problems can be reformulated as finding the ground state of the Ising model. Existing Ising solvers are mostly inspired by simulated annealing. Although annealing techniques offer scalability, they lack convergence guarantees and are sensitive to the cooling schedule. We propose solving the Ising problem by relaxing the binary spins to continuous variables and introducing an attraction potential that steers the solution toward binary spin configurations. A key property of this potential is that its combination with the Ising energy produces a Hamiltonian that can be written as a difference of convex polynomials. This enables us to design efficient iterative algorithms that require a single matrix-vector multiplication per iteration and provide convergence guarantees. We implement our Ising solver on a wide range of GPU platforms, from edge devices to high-performance computing clusters, and demonstrate that it consistently outperforms existing solvers across problem sizes ranging from small ($10^{3}$ spins) to ultra-large ($10^{8}$ spins).
@@ 1 Introduction
Combinatorial optimization problems arise in numerous fields, including social networks [yuan2018], finance [roman2019], cryptography [wang2020prime], scheduling [rieffel2014], electronic circuit design [barahona1988circuit], and biosciences [kell2012, robert2021]. The challenge with such problems is that the number of possible solutions grows exponentially with the number of variables, rendering classical algorithms slow or impractical for large-scale problems. To address this, researchers are increasingly exploring unconventional computing paradigms. A particularly promising direction is reformulating the problem as finding the ground state of a physical system, such as the Ising model. Indeed, many Nondeterministic Polynomial-time (NP)-hard problems can be naturally expressed in the Ising form, while many others can be efficiently reduced to this framework [lucas2014]. Significant effort has been directed toward developing specialized hardware and algorithms, collectively known as Ising machines, that can efficiently compute the ground states of Ising models [tanahashi2019]. As industrial optimization problems grow in complexity, building large-scale Ising solvers has become more critical than ever. Beyond its role in combinatorial optimization, finding the ground state of the Ising model is a fundamental problem in physics, offering insights into phase transitions, magnetism, and critical behaviour in condensed matter systems.
The Ising model was originally introduced by Lenz and Ising as a mathematical model for ferromagnetism [ising1925]. In this model, the dimensionless energy function is given by
@@
where $(s_{1},\ldots,s_{n})$ is the spin vector with each spin $s_{i}\in\{-1,+1\}$ (up/down spins), $J_{ij}$ is the coupling coefficient between spins $s_{i}$ and $s_{j}$ with $J_{ii}=0$ (no self-coupling), and $h_{i}$ is the external field acting on spin $s_{i}$. Solving an Ising model refers to finding a spin assignment $s^{*}_{1},\dots,s^{*}_{n}\in\{-1,1\}$ that minimizes the energy over all possible spin configurations, that is,
@@
We refer to $(s^{*}_{1},\dots,s^{*}_{n})$ as the ground state spin vector (or simply the ground state) of the Ising model.
@@
Solving large Ising models by brute force is computationally challenging, with no known polynomial-time algorithm. In fact, the problem is known to be NP-hard [barahona1982]. Consequently, rather than attempting to compute the exact ground state, practical approaches typically focus on finding good approximations to it. Over the years, many algorithmic heuristics and specialized hardware have been developed to address this challenge. We refer the reader to [mohseni2022, moy2022coupled, lo2023ising, maher2024cmos] for a survey of past and recent advances.
@@
The most widely used technique for solving Ising models is simulated annealing (SA) [kirkpatrick1983, isakov2015, goto2018boltzmann, bunyk2014], a method inspired by quantum annealing [kadow1998, giues2002, das2008]. SA uses Monte Carlo sampling guided by the Boltzmann distribution to search for the ground state. However, it relies on a heuristic cooling schedule to explore the energy landscape and does not come with any mathematical guarantees. Moreover, a gradual temperature reduction is used, which can result in slow convergence [Yavorsky2019], making SA less suited for solving large-scale problems. Variants such as noisy mean field annealing (NMFA) [King2018, Veszeli2021] and mean field annealing from a random state (MARS) [Yavorsky2019] aim to speed up convergence. However, they are typically prone to approximation errors and sensitive to hyperparameter settings. State-of-the-art techniques based on Hamiltonian dynamics, such as Simulated Bifurcation Machine (SBM) [puri2017, goto2019, goto2021, kanao2022, wang2023] and its ballistic variant (bSB) [goto2021], offer parallelism but remain sensitive to parameter tuning and cooling schedules. On the other hand, Coherent Ising Machines (CIMs) [inagaki2016, tiunov2019, honjo2021], which leverage optical hardware, suffer from issues like noise, decoherence, and the need for complex, precisely controlled hardware [yamamoto2017, pramanik2023]. More recently, neural network-based approaches have shown promise [zhang2018rbm, niazi2024, jiang2024]. However, they face inherent challenges, including limited scalability and the need for careful hyperparameter tuning [jiang2024]. Thus, there is a need to develop Ising solvers that can deliver both high-speed performance and reliable solution quality, particularly for ultra-large-scale problems.
@@
Figure 1: Overview of our Ising solver for a $2$-spin system. We relax the $2^{2}=4$ discrete energy values $(\mathcal{E}_{0},\mathcal{E}_{1},\mathcal{E}_{2},\mathcal{E}_{3})$ to a smooth energy landscape $\mathcal{E}(x_{1},x_{2})$ defined over the continuous space $(x_{1},x_{2})\in\mathbb{R}^{2}$. This energy is combined with an attraction potential $\mathcal{A}(x_{1},x_{2})$ to form the Hamiltonian $\mathcal{H}=\mathcal{E}+\mathcal{A}$. The local minima of this Hamiltonian, denoted by green dots $(\mathcal{H}_{0},\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3})$, serve as candidate ground states. With suitable parameter tuning, the minimizer $(x^{*}_{1},x^{*}_{2})$ of the Hamiltonian produces a binary vector $\mathrm{sign}(x^{*}_{1},x^{*}_{2})$ that closely matches the true ground state $\boldsymbol{s}^{*}=(s^{*}_{1},s^{*}_{2})$ of the original Ising energy $\mathcal{E}$.
@@
In this work, we propose a continuous optimization-based Ising solver. Unlike the classical Goemans-Williamson Semidefinite Program (GW-SDP) [goemans1995], our method avoids the computational bottleneck associated with large semidefinite programs (Supplementary Note 5). To improve the scalability, we drop the binary spin constraints entirely and instead couple an attraction potential with the Ising energy, resulting in a Hamiltonian with soft spin constraints. A key insight is that this Hamiltonian can be expressed as a difference of convex polynomials, enabling the use of difference-of-convex programming (DCP) [abbaszadehpeivasti2023, phan2018], which, to our knowledge, has not been applied to the Ising model. Specifically, we present two DCP-based iterative solvers that require a single matrix-vector multiplication per iteration and have just two tunable parameters. By exploiting the mathematical properties of our Hamiltonian, we prove that one of our solvers converges to a critical point, and under mild assumptions, to a local minimum, a guarantee that is difficult to obtain in general nonconvex optimization. Our solvers are well-suited for parallel execution on advanced graphical processing units (GPUs) and run efficiently across a wide range of platforms, from low-power edge devices to high-performance computing clusters. We demonstrate that they consistently outperform existing Ising solvers across a broad range of problem sizes, including ultra-large-scale instances with up to $10^{8}$ spins. Notably, we can solve a fully connected $10^{7}$-spin Ising model with nearly $50$ trillion coupling terms in just $11$ hours using four NVIDIA H100 GPUs. A quick comparison with existing Ising solvers is provided in Table 1.
@@ 2 Results
We develop a continuous-energy formulation by embedding the binary spin vector in $\mathbb{R}^{n}$ (Figure 1). As a first step, we encode the coupling coefficients in an $(n\times n)$ symmetric matrix $\mathbf{J}$ with $J_{ii}=0$, and express the homogeneous Ising problem as
$\underset{\boldsymbol{s}\in\{-1,1\}^{n}}{\min}\ \mathcal{E}(\boldsymbol{s})=-\frac{1}{2}\boldsymbol{s}^{\top}\mathbf{J}\boldsymbol{s}.$ (4)
We can identify the optimization domain $\{-1,1\}^{n}$ with the vertices of the unit hypercube in $\mathbb{R}^{n}$. We then leverage the quadratic structure of $\mathcal{E}$ to expand the domain to $\mathbb{D}=\big\{\{-\lambda,\lambda\}^{n}:\ \lambda>0\big\}$, and consider the optimization problem
@@
where $\boldsymbol{x}=(x_{1},\ldots,x_{n})$ is the new optimization variable. The domain $\mathbb{D}$ consists of radial lines emanating from the origin along the vertices of $\{-1,1\}^{n}$. Though it is larger than the domain in (4), we can compute the ground state of (4) from the global minimizers of (5). Namely, if $\boldsymbol{x}^{*}$ is a global minimizer of (5), then $\boldsymbol{s}^{*}=\mathrm{sign}(\boldsymbol{x}^{*})$ is a ground state of (4).
The formulation (5) takes us a step closer to the continuum. However, from an optimization perspective, the difficulty is that the domain $\mathbb{D}$ is a non-convex set that is not even connected. A natural idea is to consider its convex hull. However, the convex hull is the entire space $\mathbb{R}^{n}$, making the relaxation too loose and, therefore, ineffective for optimization. To address this, we introduce the attraction function that biases the minimizers towards $\mathbb{D}$. Specifically, for any fixed $\alpha,\beta>0$, we consider the attractor
@@
The parameters $\alpha$ and $\beta$ are used to control the shape of the attractor (Figure 2). From this point onward, we use $\boldsymbol{x}\in\mathbb{R}^{n}$ to denote the continuous relaxation variable.
Figure 2: Shape of the attractor. (a) Plot of the one-variable attractor for different $\alpha,\beta$ values. The global minimizers at $\pm(\alpha/\beta)^{1/2}$ are shown. (b) Plot of the attractor in two variables for $\alpha=\beta=1$, showing the four global minimizers at (1,1),(1,1),(1,1)(1,1),(1,-1),(-1,1) and (1,1)(-1,-1).
This specific form of the attractor is motivated by the Hamiltonian of nonlinear parametric oscillators [goto2018] (Supplementary Note 4). Although this provides a strong physical motivation, our primary interest in (6) is because of its mathematical properties. In particular, we can show that the global minimizers of $\mathcal{A}$ are exactly the vertices of the hypercube $\{-\lambda,\lambda\}^{n}$, where $\lambda=(\alpha/\beta)^{1/2}$ (Supplementary Note 2). As this hypercube sits inside $\mathbb{D}$ for any choice of $\lambda$, we can use this property to bias the minimizers of the Hamiltonian $\mathcal{H}=\mathcal{A}+\mathcal{E}$ toward $\mathbb{D}$. Specifically, we consider the optimization problem
@@
In summary, we begin with the exact Ising model in (4) with $\boldsymbol{s}\in\{-1,1\}^{n}$, and introduce the relaxed variable $\boldsymbol{x}\in\mathbb{D}$ in (5) to obtain an equivalent formulation. We then enlarge the domain to $\boldsymbol{x}\in\mathbb{R}^{n}$ and incorporate the attractor in (6), which leads to the unconstrained energy formulation in (7).
Unlike the original Ising problem (4), which has a finite search space, a difficulty with continuous optimization problems is that they may not have a minimizer, i.e., the problem may not be well-posed. Nevertheless, we can prove that $\mathcal{H}$ is bounded below and always has a global minimizer $\boldsymbol{x}^{*}$ (Theorem S1 in Supplementary Note 2).
Moreover, if it so happens that $\boldsymbol{x}^{*}\in\mathbb{D}$, then $\mathrm{sign}(\boldsymbol{x}^{*})$ is guaranteed to be the ground state of the original Ising problem (Supplementary Note 2). In practice, the recovered spin configuration $\mathrm{sign}(\boldsymbol{x}^{*})$ may not always coincide with a true ground state. However, by appropriately tuning the parameters $\alpha$ and $\beta$, one can steer $\boldsymbol{x}^{*}$ toward $\mathbb{D}$. We empirically demonstrate later that this typically leads to solutions of higher quality than those obtained using existing Ising solvers.
@@
A direct way of solving the optimization problem (7) is to look for a critical point of $\mathcal{H}$ by setting its gradient to zero. This leads to a system of cubic equations, which is difficult to solve in practice. Iterative schemes such as gradient descent provide a more feasible alternative, but they require evaluating the gradient $\nabla\mathcal{H}$ at every step and rely on careful step-size tuning.
We propose an iterative algorithm that has a significantly lower per-iteration cost than gradient descent and comes with a formal convergence guarantee. This is based on the observation that we can write the Hamiltonian $\mathcal{H}$ as the difference of convex functions for proper settings of $\alpha$ and $\beta$. More precisely, combining the quadratic components of $\mathcal{A}$ and $\mathcal{E}$, we have
@@
The function $f$ is convex for any $\beta>0$. On the other hand, $g$ is convex if and only if $\lambda_{\min}(\mathbf{J}+\alpha\mathbf{I})\geqslant 0$, where $\lambda_{\min}$ is the smallest eigenvalue. This can be guaranteed by making $\alpha$ sufficiently large (Supplementary Note 2). Under these conditions, the optimization problem (7) becomes
@@
where $f$ and $g$ are convex functions. This falls under the purview of difference-of-convex programming (DCP) [abbaszadehpeivasti2023, phan2018]. We consider a simple DCP algorithm called DCA [an2005], along with its accelerated variant [phan2018].
DCA builds upon the classical principle of bound optimization [dempster1977maximum]. Starting with an initialization $\boldsymbol{x}^{(0)}$, DCA generates a sequence of estimates $\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)},\cdots$ that is expected to converge to a minimizer of $\mathcal{H}$. The core idea is to exploit the convexity of $g$ to construct a global convex upper bound on fgf-g around the current estimate $\boldsymbol{x}^{(k)}$. This surrogate is then minimized to obtain the next iterate $\boldsymbol{x}^{(k+1)}$. Applied to problem (9), we obtain a simple update rule:
@@
where $\mathcal{T}$ is a linear transform followed by a pointwise nonlinearity (see Methods). In other words, the algorithm reduces to the repeated application of the operator $\mathcal{T}$, referred to as a fixed-point algorithm. The cost per iteration is just a matrix-vector multiplication. A simple example illustrating the behaviour of this iterative algorithm is shown in Figure 3. We use a two-spin Ising problem to visualize the update steps and the convergence of the iterates.
Figure 3: Optimization trajectory for a $2$-spin system. We consider the anti-ferromagnetic model $J_{12}=J_{21}=-1$, whose ground states are (1,1)(1,-1) and (1,1)(-1,1). Top row: Surface plots of the attractor $\mathcal{A}$, Ising energy $\mathcal{E}$, and the Hamiltonian $\mathcal{H}=\mathcal{A}+\mathcal{E}$. Bottom row: Corresponding contour plots for $\alpha=1$ and $\beta=2$. In the bottom-right plot, the white curves represent the optimization trajectories of our Ising solver, starting from various initial points (indicated by green dots). Depending on the initialization, the iterates converge to one of the two degenerate ground states (red dots) after five iterations.
@@
We also consider a variant of DCA that achieves improved convergence rates using Nesterov acceleration [Nesterov2018]. We refer to this as Accelerated DCA (ADCA), which is described in detail in the Methods section. The solution quality achieved by ADCA consistently outperforms DCA as the problem size increases. Although the ADCA iterations consistently converge in practice, establishing a rigorous convergence guarantee is a challenging task.
@@
where $\eta$ is a tuning parameter. It can be shown that the smallest eigenvalue of $\mathbf{J}$ is negative, so that $\lambda_{\max}(-\mathbf{J})>0$ in (10). The above choice of $\alpha$ guarantees that $\mathbf{J}+\alpha\mathbf{I}$ is positive semidefinite, ensuring the convexity of $g$ in (9). On the other hand, the choice of $\beta$ in (10) ensures boundedness of the DCA iterates for any $\alpha$ (Supplementary Note 3). We have found that this particular setting consistently yields good results.
For large models ($n\geqslant 10^{4}$), computing $\lambda_{\max}(\mathbf{J})$ can be computationally demanding. To address this, we approximate $\lambda_{\max}(\mathbf{J})$ using Wigner’s semicircle law [mehta2004rmt], following the approach in [goto2019, goto2021] for estimating eigenvalues of large matrices. Specifically, we estimate $\lambda_{\max}(\mathbf{J})$ with $2\langle\mathbf{J}\rangle\sqrt{n}$, where $\langle\mathbf{J}\rangle$ denotes the sample variance of the entries of $\mathbf{J}$:
@@
We tune the parameter $\eta$ in (10) by running a small number of iterations of our Ising solver, observing its behaviour, and adjusting $\alpha$ accordingly (Supplementary Note 6). This straightforward tuning procedure, combined with the low per-iteration cost, makes our DCA-based solvers particularly well-suited for ultra-large-scale Ising problems.
Our overall approach is summarized in Figure 1. We begin by relaxing the binary spin constraints, incorporating the attractor with the Ising energy, and expressing the resulting Hamiltonian in a DCP form. We then tune the parameters $\alpha$ and $\beta$, and apply DCA to optimize the tuned Hamiltonian. We refer to this as DOCH (Difference-Of-Convex-Hamiltonian), and its accelerated variant as ADOCH.